OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No Solution

OpenStudy (anonymous):

first bring the absolute value term on one side and everything else on the other can you do that?

OpenStudy (anonymous):

x = −1 and x = −5

OpenStudy (anonymous):

dont really get this problem

OpenStudy (anonymous):

in future use this http://www.wolframalpha.com/

OpenStudy (anonymous):

@supanick14 are you familiar with absolute value?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ya but a little rusty

OpenStudy (jdoe0001):

well one thing you'd need to keep in mind with absolute values is that when it's written −3|2x + 6| = −12 they really mean |2x+6| => +1(2x+6) AND -1(2x+6) so, "x" will end up having 2 values, one for the 1st case, the +1, and one for the other case, the -1

OpenStudy (jdoe0001):

so, −3|2x + 6| = −12 really means -3 * +1(2x+6) = -12 AND -3 * -1(2x+6) = -12

OpenStudy (jdoe0001):

so, you solve each for "x", and you'd get the values :)